# MS-IMAP - A Multi-Scale Graph Embedding Approach for Interpretable Manifold Learning

### v1

@@ 7.1 Synthetic dataset
We assess the robustness of our method using the two moons dataset, which is a 2D manifold depicting two interleaving half-circles. We sampled $N=600$ points, and set the Gaussian noise’s standard deviation to $0.12$. While spectral based methods such as UMAP are effective under relatively “modest” noise levels, their performance deteriorates significantly in the presence of larger amounts of noise (See Figure 3). As shown in Table 1, our approach shows robustness to noise and correctly clusters most points in two moons manifold, despite the large amount of noise, and outperforms competing methods. We do further experiments on more synthetic datasets in Section 10.
@@
MS-IMAP Method 1 0.75 0.73
MS-IMAP Method 2 0.89 0.87
Table 1: Comparison of clustering performance on the Two Moons datasets.
7.2 Real Datasets
We study the performance of MS-IMAP compared to other methods for real datasets. We chose a mix of financial, biological, and image datasets: the Census dataset [12] is a financial dataset containing information about individuals extracted from the 1994 US Census; the Zilionis dataset [32] is a biological dataset containing single-cell sequencing data from different types of cells, and the Animals with Attributes (AWA) [29] dataset is an image dataset containing images of animals. More information about each dataset can be found in Section 9.
As shown in Table 2, MS-IMAP Method 2 obtains the best performance across all datasets, with a tie between UMAP and MS-IMAP Method 2 in the AWA dataset. In the Census dataset, MS-IMAP Method 2 achieves a $47\%$ increase in ARI over HeatGeo, the next best method excluding our own. In the Zilionis dataset, MS-IMAP 2 achieves a $9\%$ increase in ARI over t-SNE (we found issues running the official HeatGeo code [19] on Zilionis, so were unable to obtain results for this dataset). And in AWA, we have the ARI of MS-IMAP Method 2 is slightly better than UMAP.
Dataset Census Zilionis AWA
Method / Accuracy ARI AMI ARI AMI ARI AMI
UMAP 0.05 0.08 0.45 0.71 0.73 0.80
t-SNE 0.01 0.02 0.64 0.34 0.68 0.76
HeatGeo 0.15 0.10 N/A N/A 0.65 0.74
MS-IMAP Method 1 0.15 0.14 0.69 0.76 0.72 0.80
MS-IMAP Method 2 0.22 0.15 0.70 0.76 0.74 0.81
Table 2: Clustering results comparison using ARI and AMI the Census, Zilionis, and AWA datasets. Note HeatGeo had code execution issues on the Zilionis dataset.
8 Explicit Feature Correspondence with Laplacian Scoring
The correspondence established between the features and the embedding space becomes valuable in obtaining feature importance. One simple measure for feature importance is the Laplacian score (LS) [16], which evaluates each feature based on its correlation with the graph Laplacian eigensystem. Specifically, the importance of a feature $f_{r}\in\mathbb{R}^{N}$ is determined from the importance of its corresponding coordinate $l$ in the embedding space, $\tilde{\psi}_{\mathbf{x}_{l}}$. We calculate the Laplacian score [16] with respect to the embedding features $\tilde{\psi}_{\mathbf{x}}$ using its graph Laplacian $\mathbf{L}$ and its associated degree matrix $\mathbf{D}$. This is achieved by first removing the mean and then computing the Laplacian score using:
$L_{s}(\tilde{\psi}_{\mathbf{x}})_{l}=\frac{(\tilde{\psi}_{\mathbf{x}})_{l}^{T}\mathbf{L}(\tilde{\psi}_{\mathbf{x}})_{l}}{(\tilde{\psi}_{\mathbf{x}})_{l}^{T}\mathbf{D}(\tilde{\psi}_{\mathbf{x}})_{l}}$ (11)
Smaller scores imply that the feature $(\tilde{\psi}_{\mathbf{x}})_{l}$ has a larger projection on the subspace of the eigenvectors associated with the smallest eigenvalues, indicating higher importance with respect to the graph structure. Each coordinate $l$ in the embedding space $\tilde{\psi}_{\mathbf{x}}$, namely $(\tilde{\psi}_{\mathbf{x}})_{l}$, was constructed using a single coordinate of the original feature $f_{r}$ space. Hence, the feature importance of $(\tilde{\psi}_{\mathbf{x}})_{l}$ can be explicitly interpreted as the importance of the associated original feature.
8.1 Application to Feature importance
Figure 2: AWA Dataset Laplacian Score plotline showing in order from high to low, the importance of the original features with respect to the embeddings.
We demonstrate the effectiveness of our approach using a computer vision example where providing interpretation, including feature importance, is of significant interest. Here we show in Figure (2) the estimated Laplacian score of the 85 semantic attributes for the AWA datasets sorted from high to low. Because the Laplacian score is a function of explained variance, one can argue that this score is a measure of feature information and we observe that the number of features capturing the highest scores is small (on the order of 10). This result highlights that most of the information in the data-set which correlates with the embedding is captured by a small group of variables which suggests their relatively higher importance. The Laplacian score ranking allows us to select a small number of variables for further analysis in order to better understand the output of statistical studies performed using the embedded data.
8.2 Runtime and Computational Complexity
The execution time of our method with Python code implementation with 32 cores Intel Xeon 8259CL running at 2.50Ghz and 256GB of RAM on the Cancer QC data-set of 48,969 samples and 306 features took 12.44 mins using Method 1. The computational complexity of Multi-Scale IMAP is of $O(NDlog(N))$. Additional details are provided in the appendix.
8.3 Conclusions and Limitations
Identifying the key drivers of high-dimensional datasets is a necessary ingredient to take advantage of unlabelled data in many practical applications in fields such as finance and biology. In this article, we introduce a novel contrastive learning framework for manifold learning via graph embeddings, capitalizing on both low and high-frequency information. In particular, we use SGW to construct a multi-scale graph representation of the underlying input feature space. We then use an SGD-based optimization scheme together with innovative strategies such as 3D Tensor encoding to derive the embeddings. We study the theoretical properties of the spectral graph wavelet (SGW) representation by considering functions in Paley-Wiener spaces on combinatorial graphs and prove that the SGW operator provides a more effective representation using the concept of the $\Lambda$set. Finally, we show that the embeddings are interpretable using a simple derivation of feature importance of the embedding and original feature spaces. The construction of our methodology for generating the embeddings implicitly gives us a way to tie the original and transformed feature spaces together which is lacking in current non-linear manifold learning techniques to the best of our knowledge.
Limitations: While our work has demonstrated several significant properties with strong performance on challenging datasets, it is not without limitations. Firstly, we rely on the assumption that the features, and thus the similarity measure, adequately capture information to reveal geodesic distances between manifold points. Secondly, utilizing the current optimization framework with SGD may present difficulties when extending to out-of-sample problems. Furthermore, although our embedding feature correspondence proves valuable in assessing feature importance through a global method, there remains potential for enhancement through the incorporation of local methods.
Appendix
9 Dataset Details
In Figure 3, we show example clusterings for each of the methods mentioned in Section 7.
Figure 3: Clustering results using two moons
10 Experimental Results on More Datasets
Here we show further experiments on more datasets.
Dense cluster inside a sparse circle: In this example, we sampled $500$ points from a dense cluster (using a uniform distribution) situated in the interior of a sparse cluster (using 100 points), maintaining a ratio of 5:1. UMAP relies on a well-initialized kNN graph, however the disparate densities in the two manifolds lead to erroneous edge weights, resulting in errors in segmentation. Despite utilizing the same initial graph construction as UMAP, our method, employing a multi-scale representation, is more successful in segmenting the two manifolds accurately in the presence of differing densities. HeatGeo, a competitive state-of-the-art method, does the best here, and is able to separate the dense-sparse circles from each other.
Data Dense-sparse clusters
Method/Accuracy ARI AMI
UMAP 0.75 0.72
t-SNE 0.30 0.33
ISOMAP 0.74 0.65
Diffusion Maps 0.81 0.76
HeatGeo 1.0 1.0
Ms-IMAP Method 1 0.82 0.8
Ms-IMAP Tensor-encoding Method 2 0.77 0.73
Table 3: Comparison of clustering performance on Two Moons and Dense-sparse clusters datasets.
11 HeatGeo Hyperparameter Tuning Details
@@
We tune HeatGeo on each dataset in Section 7, by tuning over the parameter space:
@@ 12 Ablation Studies of Hyperparameters for MS-IMAP
Table 5: Ablation study on MS-IMAP with method 2, varying the number of neighbors. The number of filters is kept at 5.
@@
Table 6: Ablation study on MS-IMAP with method 2, varying the number of filters. The number of neighbors is kept at 15.
@@ 15 Runtime and Computational Complexity
We have performed experimental runtime studies on empirical datasets. Experiments were performed on Virtual Server containers with 32 cores Intel Xeon 8259CL running at 2.50Ghz and 256GB of RAM. On the AWA dataset of $5,685$ samples and $2,048$ features, Method 1 and the 3D based Tensor Method 2 both run in 1.35 mins. Finally on the Cancer QC data-set of 48,969 samples and 306 features, Method 1 took 12.44 mins. The computational complexity of Multi-Scale UMAP is of $O(NDlog(N))$ for construction of the multi-scale representations which includes the $k$ nearest neighbor graph using k-d tree, the SGW transform which is $O(N)$ for each dimension of the manifold for sparse graphs. The optimization stage has a complexity which scale with the number of edges in the graph, which has a complexity of $O(kDN)$.
@@
In this work, we characterize the theoretical properties of the representation power of the SGW operator by considering functions sampled from the Paley-Wiener spaces [22] on combinatorial graphs. The Paley-Wiener spaces were introduced on combinatorial graphs in [22] and a corresponding sampling theory was developed which resembles the classical one. Pesenson proved in [22] that Paley-Wiener functions of low type are uniquely determined by their values on certain subgraphs (which are composed from set of nodes known as the uniqueness sets) and can be reconstructed from such sets in a stable way. We demonstrate that the SGW operator can represent functions $f$ that reside in the Paley-Wiener space on combinatorial graphs more efficiently than the graph Laplacian operator $\mathcal{L}$. The effectiveness of the SGW operator representation in this case can be understood in several ways. In one way, by the ability of the SGW operator to accurately represent functions with larger bandwidth, i.e., $f\in PW_{\omega^{\prime}}(G)$ where $\omega<\omega^{\prime}$.
@@
The Paley-Wiener space of $\omega$bandlimited signals is defined as follows:
@@
We summarize the main notions and definitions. We consider simple undirected unweighted and connected graph $G=(V,\mathbf{W})$, where $V$ is its set of $N$ vertices and $\mathbf{W}$ is its set of edges. The degree of $v$ is number of vertices adjacent to a vertex $v$ is and is denoted by $d(v)$. We assume that degrees of all vertices are bounded by the maximum degree denoted as
@@
A set of vertices $U\subset V$ is a uniqueness set for a space $PW_{\omega}(G)$ if we have two functions from $PW_{\omega}(G)$ that coincide on $U$ then they coincide on $V$.
Next the following definition of $L_{2}(S)$ is provided as the set of finite energy signals whose support is contained in $S$. The space $L_{2}(G)$ is the Hilbert space of all complex-valued functions $f:V\rightarrow\mathbb{C}$ with the following inner product $\left\langle f,g\right\rangle=\sum_{v\in G}f(v)\overline{g(v)}$ and the norm
$||f||=\left(\sum_{v\in V}{|f(v)|^{2}}\right)^{1/2}$ (15)
@@
[22] We say that a set of vertices $S\subset V$ is a $\Lambda$set if for any $\varphi\in L_{2}(S)$ it admits a Poincare inequality with a constant $\Lambda(S)>0$
@@
Definition above provides a tool to determine when bandlimited signals in Paley-Wiener spaces $PW_{\omega}(G)$ can be uniquely represented from their samples on a given set. The role of $\Lambda$-sets was explained and proved in the following Theorem by Pesenson [22], that shows that if $S\subset V$ then any signal $f\in PW_{\omega}(G)$ can be uniquely represented by its samples in the complement set $U=V(G)\setminus S$:
@@
Remark: Note that non-trivial uniqueness sets can not exist for functions from any Paley-Wiener subspace $PW_{\omega}(G)$ with any $\lambda_{0}\leq\omega\leq\lambda_{N}$, but can they can exist for some range $\lambda_{0}\leq\omega<\Omega$, as was shown in [22].
@@
Let $G=(V,\mathbf{W})$ be a connected graph with $N$ vertices. Assume that there exist a set of vertices $S\subset V$ for which the conditions (1)-(2) in Lemma 1 hold true. Let $\psi$ be the SGW operator using a polynomial $p(\mathbf{\cal L})$ with the coefficients $\left\{a_{k}\right\}_{k=0}^{K}$ such that $\psi_{f}=\sum_{k=0}^{K}a_{k}\mathbf{\cal L}^{k}f$. Then, for any $\varphi\in L_{2}({S})$, we have that the following inequality holds:
@@
[22] Given a connected graph $G=(V,\mathbf{W})$ graph with $N$ vertices for which the following conditions hold true: (1) For every $s\in S$ there exists $u\in U$ that is a neighbor of $s$, i.e., $w(u,s)>0$.
(2) for every $s\in S$ there exists at least one neighbor node $u\in U$ whose adjacency set intersects $S$ only over $s$.
@@
Assuming $f\in PW_{\omega}(G)$, $f$ can be efficiently represented using a polynomial $p(\mathcal{L})$ instead of a kernel function $g$, where the wavelet at node $i$ and a fixed scale $s$ is calculated using the polynomial $p(\mathbf{\cal L})$ of the Laplacian:
$\psi(i)=\sum_{k=0}^{K}a_{k}\mathbf{\cal L}^{k}\delta_{i}$ (19)
and we can write the SGW coefficients with respect to the function $\varphi$ as:
$\psi_{{\varphi}}=\sum_{l=0}^{N-1}\sum_{k=0}^{K}a_{k}\mathbf{\cal L}^{k}\varphi$
Following the assumptions of Theorem 1 implies that there exists a subset $U^{*}\subset U$ such that for every $s\in S$ there exists at least one point $u^{*}\in U^{*}$ whose adjacency set intersects $S$ only over $s$ and from this property we have
$\mathbf{\cal L}\varphi(u^{*})=-\frac{\varphi(s)}{\sqrt{d(s)d(u^{*})}},u^{*}\in U^{*},s\in S$ (20)
We will use this property to show that the set $S$ is a $\Lambda_{\psi}$-set with respect to the operator $\psi$. Using 20 and taking taking powers of the Laplacian operator considering $u^{*}\in U^{*},s\in S$ we have that:
$\mathbf{\cal L}^{k}\varphi(u^{*})=(-1)^{k}\frac{\varphi(s)}{{(d(s)d(u^{*}))^{k/2}}},u^{*}\in U^{*},s\in S$ (21)
and
$|\mathbf{\cal L}^{k}\varphi(u^{*})|\geq\frac{|\varphi(s)|}{\Lambda^{k}},u^{*}\in U^{*},s\in S$ (22)
Thus we have that $\forall\varphi\in L_{2}(S)$
$||\psi_{\varphi}||=\left(\sum_{i,j}\sum_{k=0}^{K}|a_{k}({\cal L}^{k}\varphi)_{ij}|^{2}\right)^{1/2}\geq\left(\sum_{s\in S}\sum_{k=0}^{K}|a_{k}({\cal L}^{k}\varphi(u^{*}))|^{2}\right)^{1/2}$ (23)
using similar argument as in Lemma 1 we have that
$\left(\sum_{k=0}^{K}\sum_{s\in S}|a_{k}{\cal L}^{k}\varphi(u^{*})|^{2}\right)^{1/2}=\left(\sum_{k=0}^{K}|a_{k}|^{2}\sum_{s\in S}|{\cal L}^{k}\varphi(u^{*})|^{2}\right)^{1/2}\geq||\varphi||\left(\sum_{k=0}^{K}|a_{k}|^{2}\frac{1}{\Lambda^{2k}}\right)^{1/2}$ (24)
which proves the claim of the Theorem. $\square$
Remark 1: An important property which can be observed from Theorem 3 is the following: given the SGW $\psi_{{f}}$ as a spectral representation operator, we can choose coefficients $\left\{a_{k}\right\}_{k=0}^{K}$ such that we obtain a $\Lambda_{\psi}$ set associated with the operator $\psi$, which is smaller than the $\Lambda$ set associated with the Laplacian operator $\cal L$. This implies that the operator $\psi$ provides more flexibility and better control over smoothness properties in comparison to the Laplacian operator.
@@
Remark 2: We note that the results of Theorem 4 concerning the uniqueness set are independent from the stability properties of the representation. In order to achieve stability which is important for reconstruction, it is required to construct a wavelet operator using multiple scales $t_{j},j=1,...K$, as proposed in [15], which will yield a collection of $nK$ wavelets. More specifically, we can express the function $\varphi$ using multiple scales $t_{j}$ (here we replace the previous notation of scale $s_{j}$ with $t_{j},j=1,..K$, not to confuse with nodes $s\in S$). Then for a fixed scale $t_{j}$ we have that the SGW is given by $||\psi_{\varphi}(t_{j})||=\left(\sum_{i=1}^{N}\sum_{k=0}^{K}|a_{t_{j},k}({\cal L}^{k}\varphi(i))|^{2}\right)^{1/2}$ and

### v4

@@ 11.1 Synthetic dataset
Here we show further experiments on more data, where we assess the robustness of our method using the two moons dataset, which is a 2D manifold depicting two interleaving half-circles. We sampled $N=600$ points, and set the Gaussian noise standard deviation to $0.12$. While spectral based methods such as UMAP are effective under relatively “modest” noise levels, their performance deteriorates in the presence of larger amounts of noise. As shown in Table 1 under the Moons column, our approach is robust and correctly clusters most points despite the large noise level, and competitive with the competing methods.
Data Dense-sparse clusters
@@
MS-IMAP 0.89 0.87
Table 2: Comparison of clustering performance on the two moons dataset.
12 Hyperparameter tuning details
@@
We tune UMAP on each dataset in the Experimental Results section, by tuning over the parameter space in Table 3.
Hyperparameter Set of values
n_neighbors 2, 10, 15, 20, 30, 50, 100
min_dist 0, 0.1, 0.5, 0.99
n_components 2, 3
Table 3: Space of parameters in which we tuned UMAP.
12.1 TSNE Hyperparameter Tuning Details
We tune TSNE on each dataset in the Experimental Results section, by tuning over the parameter space in Table 4.
Hyperparameter Set of values
perplexity 15, 30, 60
early_exaggeration 12, 24
n_components 2, 3
Table 4: Space of parameters in which we tuned TSNE.
12.2 PHATE Hyperparameter Tuning Details
We tune PHATE on each dataset in the Experimental Results section, by tuning over the parameter space in Table 5.
Hyperparameter Set of values
knn 3, 5, 10
decay (i.e. $\alpha$) 6, 10, 14, 18, 40
t auto, 50, 100, 150, 200
Table 5: Space of parameters in which we tuned PHATE.
12.3 HeatGeo Hyperparameter Tuning Details
We tune HeatGeo on each dataset in the Experimental Results section, by tuning over the parameter space in Table 6,
@@ 13 Ablation Studies of Hyperparameters for MS-IMAP
Table 7: Ablation study on MS-IMAP with encoding method 2, varying the number of neighbors. The number of filters is kept at 5.
@@
Table 8: Ablation study on MS-IMAP with encoding method 2, varying the number of filters. The number of neighbors is kept at 15.
@@ 16 Runtime and Computational Complexity
We have performed experimental runtime studies on empirical datasets. Experiments were performed on Virtual Server containers with 32 cores Intel Xeon 8259CL running at 2.50Ghz and 256GB of RAM. On the AWA dataset of $5,685$ samples and $2,048$ features, the 3D based Tensor and the simple concatenation Method both run in 1.35 mins. Finally on the Cancer QC data-set of 48,969 samples and 306 features, Method 1 took 12.44 mins. The computational complexity of Multi-Scale UMAP is of $O(NDlog(N))$ for construction of the multi-scale representations which includes the $k$ nearest neighbor graph using k-d tree, the SGW transform which is $O(N)$ for each dimension of the manifold for sparse graphs. The optimization stage has a complexity which scale with the number of edges in the graph, which has a complexity of $O(kDN)$.
@@
In this work, we characterize the theoretical properties of the representation power of the SGW operator by considering functions sampled from the Paley-Wiener spaces [27] on combinatorial graphs. The Paley-Wiener spaces were introduced on combinatorial graphs in [27] and a corresponding sampling theory was developed which resembles the classical one. Pesenson proved in [27] that Paley-Wiener functions of low type are uniquely determined by their values on certain subgraphs (which are composed from a set of nodes known as the uniqueness sets) and can be reconstructed from such sets in a stable way. We demonstrate that the SGW operator can represent functions $f$ that reside in the Paley-Wiener space on combinatorial graphs more efficiently than the graph Laplacian operator $\mathcal{L}$. The effectiveness of the SGW operator representation in this case can be understood in several ways. In one way, by the ability of the SGW operator to accurately represent functions with larger bandwidth, i.e., $f\in PW_{\omega^{\prime}}(G)$ where $\omega<\omega^{\prime}$. We first summarize the main notions and definitions. The space $L_{2}(G)$ is the Hilbert space of all complex-valued functions $f:V\rightarrow\mathbb{C}$ with the following inner product $\left\langle f,g\right\rangle=\sum_{v\in G}f(v)\overline{g(v)}$ and the norm
$||f||=\left(\sum_{v\in V}{|f(v)|^{2}}\right)^{1/2}$ (10)
@@
The Paley-Wiener space of $\omega$bandlimited signals $f\in L_{2}(G)$ is defined as follows:
@@
We consider a simple, undirected, unweighted, and connected graph $G=(V,\mathbf{W})$, where $V$ is its set of $N$ vertices and $\mathbf{W}$ is its set of edges. The degree of $v$ is number of vertices adjacent to a vertex $v$ is and is denoted by $d(v)$. We assume that degrees of all vertices are bounded by the maximum degree denoted as
@@
A set of vertices $U\subset V$ is a uniqueness set for a space $PW_{\omega}(G)$ if for every two functions from $PW_{\omega}(G)$ that coincide on $U$, then they coincide on $V$.
@@
[27] We say that a set of vertices $S\subset V$ is a $\Lambda$set if for any $\varphi\in L_{2}(S)$ it admits a Poincare inequality with a constant $\Lambda=\Lambda(S)>0$
@@
The definition above provides a tool to determine when bandlimited signals in Paley-Wiener spaces $PW_{\omega}(G)$ can be uniquely represented from their samples on a given set. The role of $\Lambda$-sets was explained and proved in the following Theorem by Pesenson [27], that shows that if $S\subset V$ then any signal $f\in PW_{\omega}(G)$ can be uniquely represented by its function values in the complement set $U=V(G)\setminus S$:
@@
Remark: Note that non-trivial uniqueness sets can not exist for functions from any Paley-Wiener subspace $PW_{\omega}(G)$ with any $\lambda_{0}\leq\omega\leq\lambda_{N}$, but they can exist for some range $\lambda_{0}\leq\omega<\Omega$, as was shown in [27].
@@
Let $G=(V,\mathbf{W})$ be a connected graph with $N$ vertices. Assume that there exist a set of vertices $S\subset V$ for which the conditions (1)-(2) in Lemma 1 below hold true. Let $\psi$ be the SGW operator using a polynomial $p(\mathbf{\cal L})$ with the coefficients $\left\{a_{k}\right\}_{k=0}^{K}$ such that $\psi_{f}=\sum_{k=0}^{K}a_{k}\mathbf{\cal L}^{k}f$. Then, for any $\varphi\in L_{2}({S})$, we have that the following inequality holds:
@@
Lemma 1 (Lemma 3.6 of [27])
Given a connected graph $G=(V,\mathbf{W})$, a set of vertices $S\subset V$, its complement $U=V\setminus S$, for which the following conditions hold true
  1. 1.
    For every $s\in S$ there exists $u\in U$ that is a neighbor of $s$, i.e., $w(u,s)>0$.
  2. 2.
    For every $s\in S$ there exists at least one neighbor node $u\in U$ whose adjacency set intersects $S$ only over $s$.
We require one more property of the Laplacian operation, Lemma 3 below. To this end, we use a lemma from [27],
Lemma 2 (Lemma 3.9 of [27])
If $S$ is a $\Lambda$-set, then for any $\varphi\in L_{2}(S)$ and all $t\geq 0$, $k=2^{l}$, $l=0,1,2,...$
$\frac{1}{\Lambda^{k}}\|\mathcal{L}^{t}\varphi\|\leq\|\mathcal{L}^{k+t}\varphi\|,\quad\varphi\in L_{2}(S)$
Now we are ready to prove a modification of Lemma 3.9 from [27]:
@@
Let $S$ be a $\Lambda$-set, $\varphi\in L_{2}(S)$, and let $m\geq 1$ be an integer. Then
$\frac{1}{\Lambda^{m}}\|\varphi\|\leq\|\mathcal{L}^{m}\varphi\|.$
Proof: Because every positive integer has binary representation, then we have for some $j_{i}\in\{0,1,2,...\},i=1,2,...,N$, that
$m=\sum_{i=1}^{N}2^{j_{i}}.$
Using Lemma 2, let $k=2^{j_{1}}$ and $t=2^{j_{2}}+\cdots+2^{j_{N}}$, so $m=k+t$. Then we have
$\|\mathcal{L}^{m}\varphi\|$ $=\|\mathcal{L}^{k+t}\varphi\|$
$\geq\frac{1}{\Lambda^{k}}\|\mathcal{L}^{t}\varphi\|$
$=\frac{1}{\Lambda^{2^{j_{1}}}}\|\mathcal{L}^{2^{j_{2}}+\cdots+2^{j_{N}}}\varphi\|$
We can now let $k_{2}=2^{j_{2}}$ and $t_{2}=2^{j_{3}}+\cdots+2^{j_{N}}$ so that,
$\|\mathcal{L}^{m}\varphi\|\geq\frac{1}{\Lambda^{2^{j_{1}}}}\|\mathcal{L}^{2^{j_{2}}+\cdots+2^{j_{N}}}\varphi\|$ $=\frac{1}{\Lambda^{2^{j_{1}}}}\|\mathcal{L}^{k_{2}+t_{2}}\varphi\|$
$\geq\frac{1}{\Lambda^{2^{j_{1}}}}\frac{1}{\Lambda^{k_{2}}}\|\mathcal{L}^{t_{2}}\varphi\|$
$=\frac{1}{\Lambda^{2^{j_{1}}+2^{j_{2}}}}\|\mathcal{L}^{2^{j_{3}}+\cdots+2^{j_{N}}}\varphi\|$
Continuing in this way, we eventually get,
$\|\mathcal{L}^{m}\varphi\|\geq\frac{1}{\Lambda^{m}}\|\varphi\|.$
completing the proof. $\square$
@@
Let $S\subset V$ be a set satisfying the conditions in Lemma 1, so $S$ is $\Lambda$-set. Then by Lemma 3, for any integer $k\geq 1$,
$\frac{1}{\Lambda^{k}}\|\varphi\|\leq\|\mathcal{L}^{k}\varphi\|.$
The SGW operator of $\varphi$ has the form,
$\psi_{\varphi}:=\sum_{i,j}\sum_{k=0}^{K}a_{k}\mathbf{\cal L}^{k}\varphi$
for some $K\geq 1$, $a_{k}\in\mathbb{R}$ are coefficients, and where the summation over $i$ and $j$ represents the summation over the nodes of the graph. Now we calculate,
$\|\psi_{\varphi}\|$ $=\left(\sum_{i,j}\sum_{k=0}^{K}\left|a_{k}(\mathcal{L}^{k}\varphi)_{ij}\right|^{2}\right)^{1/2}$
$=\left(\sum_{i,j}\sum_{k=0}^{K}a_{k}^{2}\left|(\mathcal{L}^{k}\varphi)_{ij}\right|^{2}\right)^{1/2}$
$=\left(\sum_{k=0}^{K}a_{k}^{2}\sum_{i,j}\left|(\mathcal{L}^{k}\varphi)_{ij}\right|^{2}\right)^{1/2}$
$=\left(\sum_{k=0}^{K}a_{k}^{2}\|\mathcal{L}^{k}\varphi\|^{2}\right)^{1/2}$
$\geq\left(\sum_{k=0}^{K}a_{k}^{2}\frac{1}{\Lambda^{2k}}\|\varphi\|^{2}\right)^{1/2}$
$=\|\varphi\|\left(\sum_{k=0}^{K}a_{k}^{2}\frac{1}{\Lambda^{2k}}\right)^{1/2}$
Letting
$\Lambda_{\psi}:=\frac{1}{\sqrt{\sum_{k=0}^{K}\frac{a_{k}^{2}}{\Lambda^{2k}}}}$
Then we have $\|\varphi\|\leq\Lambda_{\psi}\|\psi_{\varphi}\|$, so $S$ is also a $\Lambda_{\psi}$-set.
$\square$
From the assumptions and proof of Theorem 3, we get the immediate result:
Corollary 1
Let $S\subset V$ be a $\Lambda$-set. Then $S$ is also a $\Lambda_{\psi}$-set.
Remark 1: An important property which can be observed from Theorem 3 is the following: given the SGW $\psi_{{f}}$ as a spectral representation operator applied on $f$, we can choose coefficients $\left\{a_{k}\right\}_{k=0}^{K}$ such that we obtain a $\Lambda_{\psi}$ set associated with the operator $\psi$, which is smaller than the $\Lambda$ set associated with the Laplacian operator $\cal L$. This implies that the operator $\psi$ provides more flexibility and better control over smoothness properties in comparison to the Laplacian operator.
@@
Remark 2: We note that the results of Theorem 4 concerning the uniqueness set are independent from the stability properties of the representation. In order to achieve stability which is important for reconstruction, it is required to construct a wavelet operator using multiple scales $t_{j},j=1,...J$, as proposed in [18]. More specifically, we can express the function $\varphi$ using multiple scales $t_{j}$ (here we replace the previous notation of scale $s_{j}$ with $t_{j},j=1,..J$, not to confuse with nodes $s\in S$). Then for a fixed scale $t_{j}$ we have that the SGW is given by $||\psi_{\varphi}(t_{j})||=\left(\sum_{i=1}^{N}\sum_{k=0}^{K}|a_{t_{j},k}({\cal L}^{k}\varphi(i))|^{2}\right)^{1/2}$ and